OpenStudy (anonymous):

If the function y = x^5 is transformed to y = x^5 + 3, which statement is true about the transformed function?

OpenStudy (anonymous):

It is an even function. It is an odd function. It is neither an even nor an odd function. It is both an even and an odd function. It is symmetric with respect to the x-axis.

OpenStudy (anonymous):

y=x^5 + 3 can also be written as y = x^5 + 3x^0. There is a mix of even and odd exponents, so...

OpenStudy (anonymous):

It has to be both even and odd right

OpenStudy (anonymous):

actually it's neither. A function can't be both

OpenStudy (anonymous):

oh okay thank you so much

OpenStudy (anonymous):

yw